Find the multiplicative inverse of the
-1 × (-2/5) ...

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

-1 × (-2/5) = 2/5
Multiplicative inverse of 2/5 is 5/2

​Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Simplifying the number,
-1 × (-2/5) = 2/5
MULTIPLICATIVE INVERSE of 2/5  = 5/2
 
Hope it helps!
Thank you ✌️